OpenStudy (anonymous):

f(x) varies inversely with x and f(x) = –10 when x = 20. What is the inverse variation equation?

OpenStudy (ankit042):

f(x) = k/x -10 = k/20 solve for k

OpenStudy (gorv):

after solving put the value of K in equation

OpenStudy (anonymous):

i dont get it

OpenStudy (cggurumanjunath):